OpenStudy (anonymous):

Find the equation of the circle with center at (3, 2) and through the point (5, 4).

OpenStudy (anonymous):

\[(x-h)^2+(y-k)^2=r^2\] you have \((h,k)=(3,2)\) you need \(r^2\)

OpenStudy (anonymous):

find the square of the distance between the points \((3,2)\) and \((5,4)\) via \[r^2=(5-3)^2+(4-2)^2\]

OpenStudy (anonymous):

Thanks that helped